OpenStudy (anonymous):

The sum of a number and its reciprocal is 2.9. What is the number?

OpenStudy (king):

is it 2.9 or 29?

OpenStudy (anonymous):

x+1/x=2.9 (x^2+1)/x=2.9 x^2-2.9x+1=0 x=5/2 or 2/5

OpenStudy (king):

no x=5/2 or 0.4